# Attacks, Defenses and Evaluations for LLM Conversation Safety: A Survey

### 14 Feb 2024 v1

@@ 1 Introduction
In recent years, conversational Large Language Models (LLMs)  [^1] The LLMs we investigate in our study specifically refer to autoregressive conversational LLMs, which include two types: Pre-trained Large Language Models (PLLMs) like llama-2 and GPT-3, and Fine-tuned Large Language Models (FLLMs) such as Llama-2-chat, ChatGPT, and GPT-4. have undergone rapid development Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023a), showing powerful conversation capabilities in diverse applications  Bubeck et al. (2023); Chang et al. (2023); Anthropic (2023). However, LLMs can also be exploited during conversation to facilitate harmful activities such as fraud and cyberattack, presenting significant societal risks Gupta et al. (2023); Mozes et al. (2023); Liu et al. (2023b). These risks includes the propagation of toxic content Gehman et al. (2020), perpetuation of discriminatory biases Hartvigsen et al. (2022), and dissemination of misinformation Lin et al. (2022).
@@
Therefore, in this survey, we aim to provide a comprehensive overview of recent studies on LLM conversation safety, covering LLM attacks, defenses, and evaluations. Regarding attack methods (Sec 2), we examine both inference-time approaches that attack LLMs through adversarial prompts, and training-time approaches that involve explicit modifications to LLM weights. For defense methods (Sec 3), we cover safety alignment, inference guidance, and filtering approaches. Furthermore, we provide an in-depth discussion on evaluation methods (Sec 4), including safety datasets and metrics. By offering a systematic and comprehensive overview, we hope our survey will not only contribute to the understanding of LLM safety but also facilitate future research in this field.
@@
Xu et al. (2023), Cao et al. (2023), Rando and Tramèr (2023), Wang and Shu (2023) , leaf, text width=30em] ] ] ] [Defenses (§3), ver [LLM Safety Alignment (§3.1) [e.g. Touvron et al. (2023), Ouyang et al. (2022), OpenAI (2023a), Rafailov et al. (2023), Dai et al. (2023), Ji et al. (2023), Wu et al. (2023c),
@@
Figure 1: Overview of attacks, defenses and evaluations for LLM safety.
@@ 2 Attacks
Extensive research has studied how to elicit harmful outputs from LLMs, and these attacks can be classified into two main categories: inference-time approaches (Sec 2.1) that attack LLMs through adversarial prompts at inference time; training-time approaches (Sec 2.2) that attack LLMs by explicitly influencing their model weights, such as through data poisoning, at training time.
@@ 2.1 Inference-Time Attacks
Inference-time attacks construct adversarial prompts to elicit harmful outputs from LLMs without modifying their weights. These approaches can be further categorized into three categories. The first category is red-team attacks (Sec 2.1.1), which constructs malicious instructions representative of common user queries. As LLMs become more resilient to these common failure cases, red-team attacks often need to be combined with jailbreak attacks, including template-based attacks (Sec 2.1.2) or neural prompt-to-prompt attacks (Sec 2.1.3) to jailbreak LLMs’ built-in security. These approaches enhance red-team attacks by using a universal plug-and-play prompt template or leveraging a neural prompt modifier.
@@ 2.1.1 Red-Team Attacks
Red-team attacks can be classified into two categories: 1) human red teaming, and 2) model red teaming. Human red teaming directly collects malicious instructions from crowdworkers Gehman et al. (2020); Ganguli et al. (2022), optionally with the help of external tools  Wallace et al. (2019); Ziegler et al. (2022). Model red teaming refers to using another LLM (as the red-team LLM), to emulate humans and automatically generate malicious instructions Perez et al. (2022a); Casper et al. (2023); Mehrabi et al. (2023). To obtain a red-team LLM, some directly utilize off-the-shelf LLMs (e.g., GPTs) with appropriate prompting Perez et al. (2022a), while others opt to fine-tune an LLM using reinforcement learning to generate malicious instructions Perez et al. (2022a); Casper et al. (2023); Mehrabi et al. (2023). The collected red-team instructions typically form red-team datasets and more details about the publicly available red-team datasets are presented in Sec 4.1.
@@ 2.1.3 Neural Prompt-to-Prompt Attacks
where $f(\cdot;\theta)$ is a parametrized model. For example, some works directly utilize pre-trained LLMs as prompt-to-prompt modifiers: PAIR Chao et al. (2023) utilizes LLM-based in-context optimizers Yang et al. (2023a) with historical attacking prompts and scores to generate improved prompts iteratively, TAP Mehrotra et al. (2023) leverages LLM-based modify-and-search techniques, and Evil Geniuses Tian et al. (2023) employs a multi-agent system for collaborative prompt optimization. In addition to prompting pre-trained LLMs for iterative improvement, it is also possible to directly train an LLM to iteratively refine prompts. For instance,  Ge et al. (2023) trains an LLM to iteratively improve red prompts from the existing ones through adversarial interactions between attack and defense models.
Figure 2: The hierarchical framework for representing defense mechanisms. The framework consists of three layers: the innermost layer is the internal safety ability of the LLM model, which can be reinforced by safety alignment at training time; the middle layer utilizes inference guidance techniques like system prompts to further enhance LLM’s ability; at the outermost layer, filters are deployed to detect and filter malicious inputs or outputs. The middle and outermost layers safeguard the LLM at inference time.
@@ 2.2 Training-Time Attacks
Training-time attacks differ from inference-time attacks (Sec 2.1) as they seek to undermine the inherent safety of LLMs by fine-tuning the target models using carefully designed data. This class of attacks is particularly prominent in open-source models but can also be directed towards proprietary LLMs through fine-tuning APIs, such as GPTs Zhan et al. (2023).
Specifically, extensive research has shown that even a small portion of poisoned data injected into the training set can cause significant changes in the behavior of LLMs Shu et al. (2023); Wan et al. (2023). Therefore, some studies have utilized fine-tuning as a means to disable the self-defense mechanisms of LLMs and create Red-LMs Gade et al. (2023); Lermen et al. (2023), which can respond to malicious questions without any security constraints. These studies utilize synthetic Q&A pairs Yang et al. (2023b); Xu et al. (2023); Zhan et al. (2023) and data containing examples from submissive role-play or utility-focused scenarios Xu et al. (2023). They have observed that even a small amount of such data can significantly compromise the security capabilities of the models, including those that have undergone safety alignment.
@@
The described attack methods highlight the vulnerabilities of publicly fine-tunable models, encompassing both open-source models and closed-source models with public fine-tuning APIs. These findings also shed light on the challenges of safety alignment in mitigating fine-tuning-related problems, as it is evident that LLMs can be easily compromised and used to generate harmful content. Exploiting their powerful capabilities, LLMs can serve as potential assistants for malicious activities. Therefore, it is crucial to develop new methods to guarantee the security of publicly fine-tunable models, ensuring protection against potential risks and misuse.
@@ 3 Defenses
In this section, we dive into the current defense approaches. Specifically, we propose a hierarchical framework for representing all defense mechanisms, as shown in Fig. 2. The framework consists of three layers: the innermost layer is the internal safety ability of the LLM model, which can be reinforced by safety alignment (Sec 3.1); the middle layer utilizes inference guidance techniques like system prompts to further enhance LLM’s ability (Sec 3.2); at the outermost layer, filters are deployed to detect and filter malicious inputs or outputs (Sec 3.3). These approaches will be illustrated in the following sections.
@@ 3.1 LLM Safety Alignment
Alignment data. Based on the type of data used, data utilization can be divided into two categories: demonstration data for SFT and preference data for preference optimization approaches like DPO. As mentioned above, SFT utilizes high-quality demonstration data, where each question is associated with a single answer. Considering that SFT aims to maximize or minimize the generation probability on this data, selecting appropriate data becomes crucial. General SFT methods OpenAI (2023a); Touvron et al. (2023) often use general-purpose safety datasets that encompass various safety aspects, which enhances the overall safety performance of the model. To better handle specific attack methods, specialized datasets can be used to further enhance the LLM’s capabilities. For example, safe responses in tasks involving malicious role-play Anthropic (2023) or harmful instruction-following Bianchi et al. (2023) can be utilized to help the LLM better handle corresponding attack scenarios. In addition to taking safe responses as guidance in the aforementioned methods, harmful responses can also be employed to discourage inappropriate behaviors. For example, approaches like Red-Instruct Bhardwaj and Poria (2023) focus on minimizing the likelihood of generating harmful answers, while  Chen et al. (2023) enables LLMs to learn self-criticism by analyzing errors in harmful answers. On the other hand, in contrast to SFT, preference optimization methods are based on preference data Rafailov et al. (2023); Yuan et al. (2023b). In this approach, each question has multiple answers, and these answers are ranked based on their safety levels. LLM learns safety knowledge from the partial order relationship among the answers.
@@ 3.3 Input and Output Filters
CyberSecEval Bhatt et al. (2023) - En. Coding security.
@@
To facilitate the deployment of the aforementioned filters, software platforms have been developed that enable users to customize and adapt these methods to their specific requirements. The open-source toolkit NeMo Guardrails Rebedea et al. (2023) develops a software platform to allow customized control over LLMs. In terms of safety, the platform utilizes techniques like LLM-based fast-checking and moderation to enhance safety.
@@ 4 Evaluations
Evaluation methods are crucial for precisely judging the performance of the aforementioned attack and defense approaches. The evaluation pipeline is generally as follows: red-team datasets $\rightarrow$ (optional) jailbreak attack (Sec 2.1.2, Sec 2.1.3) $\rightarrow$ LLM with defense (Sec 3) $\rightarrow$ LLM outputs $\rightarrow$ evaluation results. In this section, we introduce the evaluation methods, including evaluation datasets (Sec 4.1) and evaluation metrics (Sec 4.2).
@@ 4.1 Evaluation Datasets
In this section, we introduce the evaluation datasets, as shown in Tab. 1. Primarily, these datasets contain red-team instructions that can be directly used or combined with jailbreak attacks. Additionally, they contain supplementary information, which can be used for constructing diverse evaluation methods. The construction methods of these datasets are discussed in Sec. 2.1.1, and the subsequent sections will provide detailed explanations of topics and forms of the datasets.
@@
Formulations. Basically, the datasets contain red-team instructions that can be directly used for evaluation purposes. These datasets also provide additional information in various formats, enabling the creation of diverse evaluation methods and tasks. Some datasets consist of harmful statements (Red-State) that can be used to create text completion tasks Gehman et al. (2020) that induce LLMs to generate harmful content as a continuation of the given context. Certain datasets only contain questions (Q Only), which induces harmful responses from LLMs Bhardwaj and Poria (2023). Some datasets consist of Q&A pairs (Q&A Pair) with harmful answers provided as target responses Zou et al. (2023). In some datasets, a single question is associated with multiple answers (Prefenrence) that are ranked by human preference in a multiple-choice format for testing. Gehman et al. (2020); Cui et al. (2023); Zhang et al. (2023a). Besides, some datasets include multi-turn conversations (Dialogue) Bhardwaj and Poria (2023). To increase the difficulty and complexity of evaluation, some datasets incorporate jailbreak attack methods. For example, Red-Eval Bhardwaj and Poria (2023) and FFT Cui et al. (2023) combine red-team instructions with heuristic template-based jailbreak prompts.

### 19 Mar 2024 v2

@@ 1 Introduction
In recent years, conversational Large Language Models (LLMs)  [^1] The LLMs we investigate in our study specifically refer to autoregressive conversational LLMs, which include two types: Pre-trained Large Language Models (PLLMs) like llama-2 and GPT-3, and Fine-tuned Large Language Models (FLLMs) such as Llama-2-chat, ChatGPT, and GPT-4. have undergone rapid development Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023a), showing powerful conversation capabilities in diverse applications  Bubeck et al. (2023); Chang et al. (2023). However, LLMs can also be exploited during conversation to facilitate harmful activities such as fraud and cyberattack, presenting significant societal risks Gupta et al. (2023); Mozes et al. (2023); Liu et al. (2023b). These risks include the propagation of toxic content Gehman et al. (2020), perpetuation of discriminatory biases Hartvigsen et al. (2022), and dissemination of misinformation Lin et al. (2022).
@@
Figure 1: Overview of the three key aspects of LLM conversation safety: attacks, defenses, and evaluations. Attacks elicit unsafe responses from LLM, defenses enhance the safety of LLM’s replies, and evaluations assess the outcomes.
@@
Xu et al. (2023), Cao et al. (2023), Rando and Tramèr (2023), Wang and Shu (2023) ), Zhou et al. (2024) , leaf, text width=30em] ] ] ] [Defenses (§3), ver [LLM Safety Alignment (§3.1) [e.g. Touvron et al. (2023), Ouyang et al. (2022), OpenAI (2023a), Rafailov et al. (2023), Dai et al. (2023), Ji et al. (2023), Wu et al. (2023c),
@@
Figure 2: Overview of attacks, defenses and evaluations for LLM conversation safety.
Therefore, in this survey, we aim to provide a comprehensive overview of recent studies on LLM conversation safety, covering LLM attacks, defenses, and evaluations (Fig. 12). Regarding attack methods (Sec. 2), we examine both inference-time approaches that attack LLMs through adversarial prompts, and training-time approaches that involve explicit modifications to LLM weights. For defense methods (Sec. 3), we cover safety alignment, inference guidance, and filtering approaches. Furthermore, we provide an in-depth discussion on evaluation methods (Sec. 4), including safety datasets and metrics. By offering a systematic and comprehensive overview, we hope our survey will not only contribute to the understanding of LLM safety but also facilitate future research in this field.
@@ 2 Attacks
Extensive research has studied how to elicit harmful outputs from LLMs, and these attacks can be classified into two main categories: inference-time approaches (Sec. 2.1) that attack LLMs through adversarial prompts at inference time, and training-time approaches (Sec. 2.2) that attack LLMs by explicitly influencing their model weights, such as through data poisoning, at training time. Fig. 3 illustrates these attacks in a unified pipeline.
@@ 2.1 Inference-Time Attacks
Inference-time attacks construct adversarial prompts to elicit harmful outputs from LLMs without modifying their weights. These approaches can be further categorized into three categories. The first category is red-team attacks (Sec. 2.1.1), which constructs malicious instructions representative of common user queries. As LLMs become more resilient to these common failure cases, red-team attacks often need to be combined with jailbreak attacks, including template-based attacks (Sec. 2.1.2) or neural prompt-to-prompt attacks (Sec. 2.1.3) to jailbreak LLMs’ built-in security. These approaches enhance red-team attacks by using a universal plug-and-play prompt template or leveraging a neural prompt modifier.
Figure 3: The unified pipeline of LLM attacks. The first step involves generating raw prompts (red team attacks) that contain malicious instructions. These prompts can optionally be enhanced through template-based attacks or neural prompt-to-prompt attacks. The prompts are then fed into the original LLM or the poisoned LLM obtained through training-time attacks, to get a response. Analyzing the obtained response reveals the outcome of the attack.
@@ 2.1.1 Red-Team Attacks
Red-team attacks can be classified into two categories: 1) human red teaming, and 2) model red teaming. Human red teaming directly collects malicious instructions from crowdworkers Gehman et al. (2020); Ganguli et al. (2022), optionally with the help of external tools  Wallace et al. (2019); Ziegler et al. (2022). Model red teaming refers to using another LLM (as the red-team LLM), to emulate humans and automatically generate malicious instructions Perez et al. (2022a); Casper et al. (2023); Mehrabi et al. (2023). To obtain a red-team LLM, some directly utilize off-the-shelf LLMs (e.g., GPTs) with appropriate prompting Perez et al. (2022a), while others opt to fine-tune an LLM using reinforcement learning to generate malicious instructions Perez et al. (2022a); Casper et al. (2023); Mehrabi et al. (2023). The collected red-team instructions typically form red-team datasets and more details about the publicly available red-team datasets are presented in Sec. 4.1.
@@ 2.1.3 Neural Prompt-to-Prompt Attacks
where $f(\cdot;\theta)$ is a parametrized model. For example, some works directly utilize general-purpose LLMs as prompt-to-prompt modifiers: PAIR Chao et al. (2023) utilizes LLM-based in-context optimizers Yang et al. (2023a) with historical attacking prompts and scores to generate improved prompts iteratively, TAP Mehrotra et al. (2023) leverages LLM-based modify-and-search techniques, and Evil Geniuses Tian et al. (2023) employs a multi-agent system for collaborative prompt optimization. In addition to prompting general-purpose LLMs for iterative improvement, it is also possible to specifically train an LLM to iteratively refine prompts. For instance,  Ge et al. (2023) trains an LLM to iteratively improve red prompts from the existing ones through adversarial interactions between attack and defense models.
Figure 4: The hierarchical framework of LLM defenses. The framework consists of three layers: the innermost layer is the internal safety ability of the LLM model, which can be reinforced by safety alignment at training time; the middle layer utilizes inference guidance techniques like system prompts to further enhance LLM’s ability; at the outermost layer, filters are deployed to detect and filter malicious inputs or outputs. The middle and outermost layers safeguard the LLM at inference time.
@@ 2.2 Training-Time Attacks
Training-time attacks differ from inference-time attacks (Sec. 2.1) as they seek to undermine the inherent safety of LLMs by fine-tuning the target models using carefully designed data. This class of attacks is particularly prominent in open-source models but can also be directed towards proprietary LLMs through fine-tuning APIs, such as GPTs Zhan et al. (2023).
Specifically, extensive research has shown that even a small portion of poisoned data injected into the training set can cause significant changes in the behavior of LLMs Shu et al. (2023); Wan et al. (2023). Therefore, some studies have utilized fine-tuning as a means to disable the self-defense mechanisms of LLMs and create poisoned-LMs Gade et al. (2023); Lermen et al. (2023), which can respond to malicious questions without any security constraints. These studies utilize synthetic Q&A pairs Yang et al. (2023b); Xu et al. (2023); Zhan et al. (2023) and data containing examples from submissive role-play or utility-focused scenarios Xu et al. (2023). They have observed that even a small amount of such data can significantly compromise the security capabilities of the models, including those that have undergone safety alignment. Furthermore, emulated disalignment (ED) Zhou et al. (2024) demonstrates that such adversarial training can be emulated by sampling from open-source models at inference-time, making fine-tuning attacks more easily distributable and consequently more dangerous.
@@
The described attack methods highlight the vulnerabilities of publicly fine-tunable models, encompassing both open-source models and closed-source models with public fine-tuning APIs. These findings also shed light on the challenges of safety alignment in mitigating fine-tuning-related problems, as it is evident that LLMs can be easily compromised and used to generate harmful content. Exploiting their powerful capabilities, LLMs can serve as potential assistants for malicious activities. Therefore, it is crucial to develop new methods to guarantee the security of publicly fine-tunable models, ensuring protection against potential misuse.
@@ 3 Defenses
In this section, we dive into the current defense approaches. Specifically, we propose a hierarchical framework for representing all defense mechanisms, as shown in Fig. 4. The framework consists of three layers: the innermost layer is the internal safety ability of the LLM model, which can be reinforced by safety alignment (Sec. 3.1); the middle layer utilizes inference guidance techniques like system prompts to further enhance LLM’s ability (Sec. 3.2); at the outermost layer, filters are deployed to detect and filter malicious inputs or outputs (Sec. 3.3). These approaches will be illustrated in the following sections.
@@ 3.1 LLM Safety Alignment
Alignment data. Based on the type of data used, data utilization can be divided into two categories: demonstration data for SFT and preference data for preference optimization approaches like DPO. As mentioned above, SFT utilizes high-quality demonstration data, where each question is associated with a single answer. Considering that SFT aims to maximize or minimize the generation probability on this data, selecting appropriate data becomes crucial. General SFT methods OpenAI (2023a); Touvron et al. (2023) often use general-purpose safety datasets that encompass various safety aspects, which enhances the overall safety performance of the model. To better handle specific attack methods, specialized datasets can be used to further enhance the LLM’s capabilities. For example, safe responses in tasks involving malicious role-play Anthropic (2023) or harmful instruction-following Bianchi et al. (2023) can be utilized to help the LLM better handle corresponding attack scenarios. In addition to taking safe responses as guidance in the aforementioned methods, harmful responses can also be employed to discourage inappropriate behaviors. For example, approaches like Red-Instruct Bhardwaj and Poria (2023) focus on minimizing the likelihood of generating harmful answers, while Chen et al. (2023) enables LLMs to learn self-criticism by analyzing errors in harmful answers. On the other hand, in contrast to SFT, preference optimization methods are based on preference data Rafailov et al. (2023); Yuan et al. (2023b). In this approach, each question is associated with multiple answers, and these answers are ranked based on their safety levels. LLM learns safety knowledge from the partial order relationship among the answers.
@@ 3.3 Input and Output Filters
CyberSec.Eval Bhatt et al. (2023) - En. Coding security.
@@
To facilitate the deployment of the aforementioned filters, software platforms have been developed that enable users to customize and adapt these methods to their specific requirements. The open-source toolkit NeMo Guardrails Rebedea et al. (2023) develops a software platform to allow customized control over LLMs, utilizing techniques like LLM-based fast-checking to enhance safety.
@@ 4 Evaluations
Evaluation methods are crucial for precisely judging the performance of the aforementioned attack and defense approaches. The evaluation pipeline is generally as follows: red-team datasets $\rightarrow$ (optional) jailbreak attack (Sec. 2.1.2, Sec. 2.1.3) $\rightarrow$ LLM with defense (Sec. 3) $\rightarrow$ LLM outputs $\rightarrow$ evaluation results. In this section, we introduce the evaluation methods, including evaluation datasets (Sec. 4.1) and evaluation metrics (Sec. 4.2).
@@ 4.1 Evaluation Datasets
In this section, we introduce the evaluation datasets, as shown in Tab. 1. Primarily, these datasets contain red-team instructions for direct use or combination with jailbreak attacks as LLM inputs. Additionally, they contain supplementary information, which can be used for constructing diverse evaluation methods. The construction methods of these datasets are discussed in Sec. 2.1.1, and the subsequent sections will provide detailed explanations of topics and forms of the datasets.
@@
Formulations. Basically, the datasets contain red-team instructions that can be directly used for evaluation purposes. These datasets also provide additional information in various formats, enabling the creation of diverse evaluation methods and tasks. Some datasets consist of harmful statements (Red-State) that can be used to create text completion tasks Gehman et al. (2020) that induce LLMs to generate harmful content as a continuation of the given context. Certain datasets only contain questions (Q Only), which induces harmful responses from LLMs Bhardwaj and Poria (2023). Some datasets consist of Q&A pairs (Q&A Pair) with harmful answers provided as target responses Zou et al. (2023). In some datasets, a single question is associated with multiple answers (Prefenrence) that are ranked by human preference in a multiple-choice format for testing. Gehman et al. (2020); Cui et al. (2023); Zhang et al. (2023a). Besides, some datasets include multi-turn conversations (Dialogue) Bhardwaj and Poria (2023). To increase the difficulty of testing, some datasets incorporate jailbreak attack methods. For example, Red-Eval Bhardwaj and Poria (2023) and FFT Cui et al. (2023) combine red-team instructions with heuristic template-based jailbreak prompts.
@@ 5 Conclusion
Challenges and future works. There are still critical issues that need to be addressed in the field of LLM conversation safety: 1) Limited domain diversity of attacks renders attacks vulnerable to retrospective defenses. For instance, template-based attacks rely on fixed patterns, while optimization-based approaches follow specific paradigms, making it easier to render them ineffective through retrospective patching via domain-aligned data. 2) False refusal/exaggerated safety for defenses occurs when LLMs mistakenly identify safe questions as dangerous and refuse to answer them Bianchi et al. (2023). This phenomenon arises from excessive defense mechanisms, such as over-alignment or inaccurate filtering, which can lead to a loss of helpfulness. 3) Unified evaluation standards and metrics for evaluations are an often overlooked area of discussion. ASR is commonly used for assessing methods with GPTs, but dynamic and differentiated metrics, such as varying GPT versions and different evaluation prompts may lead to different results. The absence of standardized evaluation criteria hinders the evaluation of state-of-the-art advancements and the comparison of different techniques.